A bullet of mass 1g shot vertically up by a gun gives a force 2N for 0.001s.Calculate the maximum height attained by a bullet. Neglect air friction and take g=10ms-2. Solve this problem…………..

Grade:11

1 Answers

vikas askiitian expert
13 years ago

force = change in momentam/time

2 = mv-mu/t

u=0 so

             v=2m/s

we have v2=u2+2as

     final velocity when particle reaches highest point is 0

     so 0 =u2-2gHmax

               Hmax =  u2/2g=4/20=0.2m ans